OpenStudy (anonymous):

|r - 8| = 5 Please help me solve this equation.

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

One possible answer is obviously 13.

OpenStudy (shamil98):

|r-8|=5 gives you two equations r - 8 = 5 and r - 8 = -5 just solve for r.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

ah okay, so r would equal 13 or 3.

OpenStudy (anonymous):

r would equal 3

OpenStudy (anonymous):

Wouldn't it also equal 13?

OpenStudy (anonymous):

crap , lol ya i gues it would , umm well its either one , i cant help farther than that sorry

OpenStudy (anonymous):

Thank you for your help.

OpenStudy (anonymous):

welcome